# Scalable Vision Language Model Training via High Quality Data Curation

### v1

@@ 6.2 Vision Language Model Pretrain
VLM pretraining benefits from higher-quality and larger-scale visual understanding data effectively. Previous works, such as BLIP2 Li et al. (2023) and LLaVA Liu et al. (2024b), pretrain the model with relatively low-quality caption datasets Li et al. (2023). With the advance in data engineering works, MiniCPM-V Yao et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c), and QwenVL Bai et al. (2023); Wang et al. (2024) series explore expanding high-quality visual understanding data sizes to improve model performance. In this work, we futher reveal model performance dynamics w.r.t. data quality and training data size, which are largely unexplored in previous works.
@@ 6.3 Visual Instruction Tuning
LLaVA Liu et al. (2024b) first defines visual instruction tuning and provides a baseline for VLM SFT data curation. Subsequent LLaVA series models Liu et al. (2023, 2024a); Li et al. (2024a) refine the visual instruction tuning datasets and achieve significantly better model performance. BLIP3 Xue et al. (2024) incorporates image-text interleaved data into visual instruction tuning, while CogVLM Hong et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c) and QwenVL series Wang et al. (2024) models explore using video question answering data for VLM SFT. In this paper, we elaborate the guidelines for the design of visual instruction datasets, providing valuable references for VLM training.
@@ 7 Conclusions
In this work, we introduce SAIL-VL-2B, a compact opensource vision language model with SOTA visual comprehension performance. We propose a scalable caption data construction pipeline, and curate SAIL-Caption, a caption dataset with large quantity and the highest quality among opensource alternatives. Pretraining on SAIL-Caption, we demonstrate that even compact VLMs can benefit from scaled up training data size. We further present data size scaling laws that SAIL-VL’s visual comprehension performance improves logarithmically as training data size increases. For visual instruction tuning, we elaborate several key guidelines for SFT data curation, guided by which we curate SFT datasets with balanced distribution. Expected data size scaling laws are observed during SFT on the curated datasets, and the phased SFT strategy further improves the logarithmic scaling curves to be near-linear. We evaluate SAIL-VL-2B on 18 opensource VLM benchmarks, and SAIL-VL outperforms existing VLMs of comparable sizes consistently either in overall performance or domain-specific abilities, depicting promising prospects in real-world applications.
@@ 8 Limitations
Acknowledgments
Our model is built upon many outstanding open-source projects, and we are grateful for their contributions. We extend special thanks to the InternVL team, Qwen team for their great base models, and to the BAAI team (Infinity-MM), Molmo team for their generous release of data.
@@ Appendix A Authorship and Credit Attribution
Hongyuan Dong: scalable data annotation pipeline, caption data construction, and data distribution evaluation
Weijie Yin: data collection and caption data construction
@@
Weijie Yin: Pretrain-alignment stage, data distribution ablation
Hongyuan Dong: Pretrain-advance stage, data distribution ablation, VLM visual understanding ability evaluation, data size scaling laws
@@
Zijiang Kang: SFT-instruction and SFT-preference, SFT data curation guidelines, phased SFT strategy and SFT data size scaling laws
Weijie Yin: SFT-knowledge and SFT-preference, phased SFT strategy and SFT data size scaling laws
@@
Weijie Yin: data collection, evaluation pipeline
@@
Xiao Liang, Chao Feng and Jiao Ran: resource allocation, project management.
Appendix B SAIL-Caption Showcase
B.1 Caption Data Quality Assessment
To evaluate the caption data quality of SAIL-Caption and opensource alternatives, we curate an evaluation subset for each recaption dataset as the test set for evaluation. Specifically, we randomly select 500 samples from DataComp and SA1B, respectively, and ask GPT4O OpenAI (2024) to generate ground truth captions for subsequent evaluation. SAIL-Captioner is tasked to annotate detailed captions for these samples, and the annotations are evaluated by GPT4O according to the precision and recall of visual elements. We set SA1B QwenVL Caption Data (2024b) and DataComp LLaVA Caption Li et al. (2024b) as baseline datasets, and their data quality is evaluated within the same suite. SA1B QwenVL Caption Data (2024b) employs a finetuned QwenVL model to annotate Chinese dense caption on the SA1B dataset. DataComp LLaVA Caption Li et al. (2024b), on the other hand, trains a customized version of LLaVA model to perform annotation. As shown in Table 6, SAIL-Caption-DataComp and SAIL-Caption-SA1B achieve significantly higher scores than previous baseline datasets, demonstrating its advantage in data quality.
Dataset Language # Sample Captioner Qualtiy Score
[…]
Table 6: Data quality evaluation results of SAIL-Caption and other opensource caption datasets. We curate a 500-sample subset for each dataset for evaluation.
B.2 SAIL-Caption Showcases
Figure 7: SAIL-Caption showcases. Image and questions are curated from SA1B Kirillov et al. (2023) and DataComp Gadre et al. (2024). Highlighted texts denote visual elements depicted by SAIL-Caption but neglected in other caption datasets.
We curate several image samples from SA1B Kirillov et al. (2023) and DataComp Gadre et al. (2024) as demonstrations to compare the quality of SAIL-Caption with existing opensource caption datasets. We compare SAIL-Caption with SA1B-QwenVL-Caption and DataComp-LLaVA-Caption, and showcases are shown in Figure 7. As the demonstrations show, SAIL-Caption encompasses more detailed visual elements than other alternative datasets in both English and Chinese, which coincides with quantified caption quality evaluation results shown in Section 3.1, underscoring its superior data quality.
@@ Appendix C SAIL-VL Showcases
Figure 8: SAIL-VL showcases. Image and questions are curated from LLaVABench Li et al. (2024a).
We task SAIL-VL to tackle vision-based language interactions in Figure 8. The images and input questions are selected from LLaVABench Li et al. (2024a). SAIL-VL demonstrates marked capabilities in language interactions and visual comprehension of various input visual element types, demonstrating powerful visual instruction following abilities.
Appendix D Visual Understanding Benchmark
@@
To inspect SAIL-VL’s visual understanding performance during pretraining stages, we curate a series of visual understanding benchmarks for evaluation. To be specific, we focus on evaluating model performance in detailed captioning and OCR tasks in both English and Chinese, which are also the optimization objectives of pretraining stages. We list the basic information of the selected visual understanding benchmarks in Table 7.
@@
Benchmarks.
DetailCaps-4870 Dong et al. (2024) encompasses images from a wide range of publicly available datasets, including COYO Byeon et al. (2022), LAION Schuhmann et al. (2021), CC Changpinyo et al. (2021), Flickr Young et al. (2014), SBU Ordonez et al. (2011), and COCO Chen et al. (2015), as well as ground truth detail captions generated by powerful VLM APIs. To ensure evaluation reliability, we task human annotators to refine this dataset by removing hallucinatory descriptions, correcting descriptive errors, and supplementing missing details. The dataset correction is fulfilled in Chinese, with GPT4O OpenAI (2024) deployed by Microsoft Azure as the translator. We adopt both the corrected Chinese captions and translated English captions for evaluation.
@@
We evaluate SAIL-VL’s caption performance on the DetailCaps-4870 benchmark with GPT evaluation. Provided with three ground truth captions and a candidate caption, GPT is tasked to score the candidate caption based on the precision and recall of the visual elements. For OCR tasks, we compute the ANLS score Biten et al. (2019) between the predicted OCR contents and the ground truth one, resulting in scores ranging from 0 to 1, with higher score indicating better prediction quality.

### v3

@@ 6.2 Vision Language Model Pretrain
VLM pretraining benefits from higher-quality and larger-scale visual understanding data effectively. Previous works, such as BLIP2 Li et al. (2023b) and LLaVA Liu et al. (2024b), pretrain the model with relatively low-quality caption datasets Li et al. (2023b). Subsequent works, such as MiniCPM-V Yao et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c), and QwenVL Bai et al. (2023); Wang et al. (2024b) series, explore expanding high-quality visual understanding data sizes to improve model performance. In this work, we further reveal model performance dynamics w.r.t. SFT data quality and size, which are largely unexplored in previous works.
@@ 6.3 Visual Instruction Tuning
LLaVA Liu et al. (2024b) first defines visual instruction tuning and provides a baseline for VLM SFT data curation. Subsequent LLaVA series models Liu et al. (2023, 2024a); Li et al. (2024a) refine the visual instruction tuning datasets and achieve significantly better model performance. BLIP3 Xue et al. (2024) incorporates image-text interleaved data into visual instruction tuning, while CogVLM Hong et al. (2024), InternVL Chen et al. (2024d); Team (2024a); Chen et al. (2024c) and QwenVL series Wang et al. (2024b) models explore using video question answering data for VLM SFT. In this paper, we elaborate the guidelines for the design of visual instruction datasets, providing valuable references for VLM training.
@@ 7 Conclusions
In this work, we introduce SAIL-VL, an opensource vision language model series with SOTA performance. We propose a scalable caption data construction pipeline and curate SAIL-Caption, a large-scale caption dataset with the highest quality among opensource alternatives. Equipped with SAIL-Caption, we conduct large-scale pretraining with up to 655B tokens and demonstrate that even compact VLMs can benefit from scaled up training data size. We further present data size scaling laws that SAIL-VL’s visual comprehension performance improves logarithmically as training data size increases. For visual instruction tuning stages, we elaborate on several key guidelines for high-quality SFT data curation, guided by which we curate our SFT-Instrcution dataset, a high-quality SFT data collection exhibiting improved model performance scaling curves than opensource alternatives during model training. The phased SFT strategy used in SAIL-VL SFT further improves the scaling curves from logarithmic to near-linear. We evaluate SAIL-VL on 18 opensource VLM benchmarks, and our model outperforms existing VLMs of comparable sizes consistently either in overall performance or domain-specific abilities, depicting promising prospects in real-world applications.
@@ 8 Limitations
We also point out that although SAIL-VL’s training process is designed carefully, models may generate hallucinated, biased, or harmful information under certain circumstances, which will be further discussed and mitigated in our future works.
@@ Appendix A Authorship and Credit Attribution
Hongyuan Dong, Weijie Yin
@@
Hongyuan Dong, Weijie Yin
@@
Zijiang Kang, Weijie Yin
@@
Weijie Yin
@@
Xiao Liang, Chao Feng, and Jiao Ran
Appendix B SAIL-VL Model Card
We provide a simplified model card for the proposed SAIL-VL-2B and SAIL-VL-8B model.
SAIL-VL-2B SAIL-VL-8B
# Parameter 1.85B 7.95B
LLM Qwen2.5-1.5B Qwen2.5-7B
Vision Encoder InternViT-300M
Resolution $448\times\{\{1\times 1\},...,\{2\times 5\}\}$
Adapter 2-layer MLP
Token Merge 2$\times$2
# Max Visual Token 2560
Table 7: SAIL-VL-2B and SAIL-VL-8B model card.
@@ Appendix C SAIL-VL Showcases
Figure 6: SAIL-VL-8B showcases. We include both English and Chinese queries with various input images.
We task SAIL-VL-8B to tackle vision-based language interactions in Figure 6. The images are selected from the internet, and input questions are set to cover common queries in real-world human interactions. SAIL-VL demonstrates marked capabilities in language interactions in both English and Chinese. It also exhibits marked visual comprehension abilities for various input visual element types. Our model recognizes famous landmarks, buildings, and artworks, demonstrating a vast reservoir of world knowledge. It is also worth noticing that SAIL-VL also performs well in meme understanding. It not only perceives the visual elements accurately, but also points out the contrast that makes the meme humorous, exhibiting powerful visual comprehension and language interaction abilities.
Appendix D SAIL-Caption
D.1 Caption Data Quality Assessment
Dataset Language Captioner GPT Eval Human Eval
DataComp-LLaVA-Caption EN LLaVA-Captioner 51.14 70.0
SAIL-Caption-DataComp EN SAIL-Captioner 61.50 87.2
SA1B-QwenVL-Caption CN QwenVL-Captioner 63.82 74.6
SAIL-Caption-SA1B CN SAIL-Captioner 71.36 88.2
BLIP3-KALE EN 2B VLM 59.88 73.2
SAIL-Caption-KALE EN SAIL-Captioner 61.50 80.6
Table 8: Data quality evaluation results of SAIL-Caption and other opensource caption datasets. The evaluation results of our SAIL-Captioner are marked with . The quality scores from GPT and human evaluation are rescaled to [0, 100] for simplicity.
Datasets.
To evaluate the caption data quality of SAIL-Caption and opensource alternatives, we curate an evaluation subset for each recaption dataset as the test set. Specifically, we randomly select 500 samples from three recaption datasets listed in Table 8. SA1B-QwenVL-Caption employs a finetuned QwenVL Data (2024a) model to annotate Chinese dense captions on the SA1B dataset. DataComp-LLaVA-Caption, on the other hand, trains a customized version of LLaVA Liu et al. (2023) model to perform annotation. BLIP3-KALE tasks CogVLM-18B Wang et al. (2023) and Mistral-8B-Instruct Jiang et al. (2024) to generate knowledge-augmented detail captions, and then distill this pipeline into a 2B VLM to annotate the DataComp-1B Gadre et al. (2024) dataset.
We instruct Azure GPT4O-20240513 OpenAI (2024) to generate ground truth captions for the curated evaluation sets. We also use our SAIL-Captioner to generate detail captions on the evaluation sets. Captions from the original recaption dataset and those generated by our SAIL-Captioner model are then compared with the ground truth ones for evaluation.
GPT Evaluation.
We first conduct GPT evaluation for efficiency. We feed the candidate captions and ground truth ones to Azure GPT4O API, and ask the model to judge the candidate caption quality based on the precision and recall of visual elements. The detailed prompt used for GPT evaluation is shown in Figure 7.
Figure 7: GPT and human evaluation prompts for SAIL-Caption and other opensource caption datasets.
Human Evaluation.
To further validate the advantage of the data quality of our SAIL-Caption dataset, we also task human experts to evaluate the data quality. We randomly select a 100-sample subset from each dataset and instruct the annotators to judge candidate caption quality based on the original image. As shown in Figure 7, human annotators are given two candidate captions from both the baseline dataset and SAIL-Caption simultaneously. The experts are required to provide quality scores for the given captions, as well as a Good-Same-Bad (GSB) judgment reflecting more fine-grained data quality differences.
In the inspection of 10% annotated samples, we observe a 95%+ accuracy, verifying the reliability of our human evaluation results.
Evaluation Results.
As shown in Table 8, SAIL-Caption-DataComp, SAIL-Caption-SA1B, and SAIL-Caption-KALE achieve significantly higher scores than previous baseline datasets in both GPT and human evaluation. These results demonstrate the leading performance of our SAIL-Captioner model and the advantage in SAIL-Caption’s data quality.
Figure 8: GPT and human evaluation prompts for SAIL-Caption and other opensource caption datasets.
We also show the GSB evaluation results in Figure 8. The GSB comparison reflects more fine-grained caption quality differences in candidate captions than a single rating. In the GSB evaluation, SAIL-Captioner achieves 87%, 91%, and 79% win+tie rates against SA1B-QwenVL-Caption, DataComp-LLaVA-Caption, and BLIP3-KALE, respectively, exhibiting marked quality advantages.
We attribute the leading performance of SAIL-Captioner to the simple-yet-effective data distillation pipeline. SAIL-Captioner develops visual understanding abilities effectively from reference data annotated by powerful VLM APIs, enabling large-scale high-quality data generation with limited resources.
D.2 SAIL-Caption Showcases
Figure 9: SAIL-Caption showcases versus SA1B-QwenVL-Caption, DataComp-LLaVA-Caption, and BLIP3-KALE. Images are curated from SA1B, DataComp and BLIP3-KALE.
We curate several image samples from SA1B Kirillov et al. (2023), DataComp Gadre et al. (2024), and BLIP3-KALE Awadalla et al. (2024) as demonstrations to compare the quality of SAIL-Caption with existing opensource caption datasets. We compare SAIL-Caption with SA1B-QwenVL-Caption, DataComp-LLaVA-Caption, and BLIP3-KALE. Showcases are shown in Figure 9. As the demonstrations show, SAIL-Caption encompasses more detailed visual elements than other alternative datasets in both English and Chinese. Observations drawn from these showcases coincide with quantified caption quality evaluation results shown in Section D.1, underscoring the leading data quality of SAIL-Caption.
A subset of the SAIL-Caption dataset with considerable data size will be released to promote opensource VLM research.
Appendix E Visual Understanding Benchmark
@@
To inspect SAIL-VL’s visual understanding performance during pretraining stages, we curate a series of visual understanding benchmarks for evaluation. To be specific, we focus on evaluating model performance in detailed captioning and OCR tasks in both English and Chinese, which are also the optimization objectives of SAIL-VL and opensource VLMs’ pretraining stages. We list the basic information of the selected visual understanding benchmarks in Table 9.
Benchmarks.
DetailCaps-4870 Dong et al. (2024) encompasses images from a wide range of publicly available datasets, including COYO Byeon et al. (2022), LAION Schuhmann et al. (2021), CC Changpinyo et al. (2021), Flickr Young et al. (2014), SBU Ordonez et al. (2011), and COCO Chen et al. (2015), as well as ground truth detail captions generated by powerful VLM APIs. We use the human-refined version of DetailCaps-4870 for evaluation and adopt both the corrected Chinese captions and translated English captions for multilingual evaluation.
@@
We evaluate SAIL-VL’s caption performance on the DetailCaps-4870 benchmark with GPT evaluation. Provided with three ground truth captions and a candidate caption, GPT is tasked to score the candidate caption based on the precision and recall of the visual elements. For OCR tasks, we compute the ANLS score Biten et al. (2019) between the predicted OCR contents and the ground truth ones, resulting in scores ranging from 0 to 1. The higher score indicates better prediction quality.
@@
Appendix F SFT Data Quality Evaluation
In this section, we show the detailed instructions for SFT data quality evaluation. As shown in Figure 10, human experts to annotate the challenging, complexity, and relevance scores for our three-stage SFT data.
Figure 10: Detailed instructions for human experts to judge SFT data quality.
Appendix G Generalizing Visual Understanding Abilities to Instruction Following Tasks
In this part, we investigate the correlation between VLM’s visual understanding and instruction following abilities. We train SAIL-VL pretrained checkpoints from the pretrain-advance stage with exponentially larger training data sizes, and train them through either LLaVA-Next Liu et al. (2024a) SFT data or our SFT-instruction data.
We plot the correlation of pretrained models’ visual understanding performance and SFT models’ opensource benchmark performance in Figure 11. A notable correlation is observed across different training strategies. As the VLM gains stronger visual understanding abilities during pretraining, its visual instruction following performance after SFT is improved accordingly, even if trained with different visual instruction tuning datasets. We quantify this correlation with Pearson correlation ($\rho$) and coefficient of determination ($R^{2}$). It turns out that SAIL-VL’s pretrained visual understanding performance and SFT visual instruction following performance share a significant correlation. In experiments with SFT-Instruction data collection, pretrained model performance and SFT model scores share a Pearson correlation coefficient $\rho=0.97$ and a coefficient of determination $R^{2}=0.94$. For LLaVA-Next SFT data experiments, we observe an even stronger correlation with $\rho=0.99$ and $R^{2}=0.98$. These correlation results illustrate the generalization of model abilities across training stages and tasks, validating the necessity of pretraining VLMs for more robust visual understanding abilities.
Appendix H Experiment Details
H.1 Experiment settings
Storage.
The training data of SAIL-VL’s pretraining and SFT stages is stored on our Hadoop file system (HDFS) for persistent storage. Training data is fetched in a stream fashion during model training, making possible training with large scale distributed data storage.
Training framework.
Training resources.
We conduct experiments with Huawei 910B x86 NPU asc . To train the SAIL-VL-2B model, we allocate 90,053 NPU hours for pretraining and 10,992 NPU hours for SFT stages, resulting in a total of 101,045 NPU hours in model training. For the SAIL-VL-8B model, we use 26M samples in pretraining for efficiency, and the same data collections as the 2B model are used in SFT stages. The 8B model consumes 19,575 NPU hours to train, where 6,672 and 12,903 NPU hours are allocated in pretraining and SFT stages, respectively.
H.2 Experiment Results.
We show model training details for both SAIL-VL-2B and SAIL-VL-8B models in Table 10 and Table 11. We add InternVL2 Team (2024a), InterVL2.5 Chen et al. (2024c), and Aquila Gu et al. (2024) series models as supplementary baselines for evaluation.
Figure 11: The correlation between SAIL-VL pretrained checkpoints’ understanding performance and their performance on opensource benchmarks after SFT. “OS BMK Score” stands for average score on opensource benchmarks in our evaluations.
Benchmark SAIL-VL Qwen2-VL InternVL2 InternVL2.5 InternVL2.5-MPO Aquila DeepSeekVL-2-Tiny
Overall Performance
Opensource Average 69.1 64.4 62.9 66.5 67.7 66.5 67.0
   General VQA 60.4 58.3 55.6 62.6 63.1 59.8 59.4
   OCR VQA 75.9 72.5 68.2 68.7 71.1 71.7 74.4
   Math&Knowledge 79.0 59.0 70.6 73.0 75.3 75.4 71.3
   Hallucination 66.2 62.9 61.7 66.2 64.5 62.9 63.6
General VQA
MMStar (2024b) 55.1 46.3 50.5 53.5 54.3 54.7 49.9
$\text{MMBench}_{\text{DEV}}$ (2024c) 72.4 68.8 70.3 73.1 72.5 74.4 68.3
$\text{MMMU}_{\text{VAL}}$ (2024) 40.1 39.9 34.2 40.7 41.2 44.1 39.6
MME (2023) 1969 1923 1859 2090 2123 1808 1910
$\text{SEEDBench}_{\text{IMG}}$ (2023a) 74.7 72.0 70.9 73.4 73.2 73.9 72.5
RealWorldQA (2024) 63.8 60.9 56.7 60.9 60.7 64.1 64.8
MMVet (2024b) 46.1 51.2 40.4 61.7 64.0 42.7 52.8
OCR VQA
[…]
$\text{DocVQA}_{\text{VAL}}$ (2021) 89.2 88.7 86.0 87.4 87.8 85.0 88.6
$\text{InfoVQA}_{\text{VAL}}$ (2022) 67.2 63.4 57.5 61.6 61.6 60.5 63.8
$\text{ChartQA}_{\text{TEST}}$ (2022) 81.0 70.6 71.7 73.3 70.9 76.6 81.2
$\text{TextVQA}_{\text{VAL}}$ (2019) 75.7 78.8 73.4 76.4 77.2 76.4 80.5
$\text{OCRVQA}_{\text{TEST}}$ (2019) 58.5 54.3 36.2 28.3 40.0 51.3 51.4
OCRBench (2024d) 806 794 786 789 846 772 808
Math&Knowledge
$\text{MathVista}_{\text{MINI}}$ (2023) 62.8 45.0 46.8 51.1 55.3 59.4 54.5
$\text{ScienceQA}_{\text{VAL}}$ (2022) 95.3 73.0 94.4 94.9 95.3 91.4 88.1
Hallucination
HallusionBench (2024) 45.7 38.3 38.2 42.5 39.2 42.1 38.4
POPE (2023) 86.7 87.6 85.3 89.9 89.8 83.6 88.8
Table 10: Complete evaluation results for SAIL-VL-2B and opensource VLMs of comparable sizes. Denotations are defined the same as Table 4.
Benchmark SAIL-VL Qwen2-VL InternVL2 InternVL2.5 InternVL2.5-MPO DeepSeekVL-2-Small
Overall Performance
Opensource Average 74.5 73.0 70.0 73.2 74.3 72.7
   General VQA 68.3 68.5 66.6 70.1 71.2 66.8
   OCR VQA 79.8 79.6 72.6 75.0 76.3 79.0
   Mah&Knowledge 83.3 71.0 78.4 81.5 83.2 79.0
   Hallucination 68.7 67.5 64.5 69.5 69.7 65.3
General VQA
MMStar (2024b) 64.2 58.3 61.6 62.5 65.3 57.7
$\text{MMBench}_{\text{DEV}}$ (2024c) 79.5 79.5 80.3 83.1 83.3 78.1
$\text{MMMU}_{\text{VAL}}$ (2024) 48.2 50.9 47.6 52.4 52.8 47.6
MME (2023) 2244 2321 2215 2339 2321 2149
$\text{SEEDBench}_{\text{IMG}}$ (2023a) 75.5 75.3 75.4 77.0 76.9 76.8
RealWorldQA (2024) 71.9 69.7 64.7 69.9 70.2 70.2
MMVet (2024b) 58.3 62.6 57.7 62.1 66.8 60.3
OCR VQA
[…]
$\text{DocVQA}_{\text{VAL}}$ (2021) 92.2 93.7 90.8 91.8 92.1 92.3
$\text{InfoVQA}_{\text{VAL}}$ (2022) 75.2 75.9 61.5 75.5 76.2 72.5
$\text{ChartQA}_{\text{TEST}}$ (2022) 84.6 81.6 82.0 82.9 77.6 84.6
$\text{TextVQA}_{\text{VAL}}$ (2019) 77.7 83.8 77.6 79.0 79.2 83.3
$\text{OCRVQA}_{\text{TEST}}$ (2019) 61.4 56.2 38.1 29.5 36.7 54.5
OCRBench (2024d) 835 833 746 819 880 834
Math&Knowledge
$\text{MathVista}_{\text{MINI}}$ (2023) 68.4 57.3 59.4 65.4 68.5 61.8
$\text{ScienceQA}_{\text{VAL}}$ (2022) 98.2 84.6 97.4 97.6 97.9 96.2
Hallucination
HallusionBench (2024) 52.2 48.5 44.6 50.1 50.3 41.2
POPE (2023) 85.2 86.5 84.4 88.8 89.1 89.4
Table 11: Complete evaluation results for SAIL-VL-8B and opensource VLMs of comparable sizes. Denotations are defined the same as Table 4.